# Towards Unpolarized GPDs from Pseudo-Distributions

### v1

@@ 1 Introduction
GPDs provide therefore an ideal setup where, at least for the time-being, only lattice QCD can provide systematically controlled information. Besides DVCS and DVMP, an increasing number of exclusive processes with enhanced sensitivity to GPDs are being studied, but their experimental exploration is only at its beginning: double DVCS Guidal and Vanderhaeghen (2003); Belitsky and Mueller (2003), where a recent study of measurability looks promising Deja et al. (2023), di-photon production Pedrak et al. (2020); Grocholski et al. (2022), photon meson pair production Boussarie et al. (2017); Duplančić et al. (2018), or the general category of single diffractive hard exclusive processes Qiu and Yu (2022, 2023a). The maturity of this new phenomenology of GPDs in the perspective of the EIC therefore complements the progress of the extraction of GPDs on the lattice.
@@
For the reader interested in the phenomenology of GPDs beyond the framework of lattice calculations, we summarize in Fig. 2 the main physics results of this study, in the form of six moments of the isovector GPDs $H^{u-d}$ and $E^{u-d}$:
$\int_{-1}^{1}\mathrm{d}x\,x^{n-1}\begin{pmatrix}H^{u-d}\\ E^{u-d}\end{pmatrix}(x,\xi,t)=\sum_{k=0\textrm{ even}}^{n-1}\begin{pmatrix}A_{n,k}(t)\\ B_{n,k}(t)\end{pmatrix}\xi^{k}\,.$ (1)
For brevity, the results are quoted at the scale $\mu=2$ GeV in the form of a dipole fit with the value at $t=0$ and a dipole mass $\Lambda_{n,k}$ following the example:
@@
More sophisticated treatment of the data is presented starting from section 5 of this document. The quoted uncertainty includes an evaluation of excited state contamination on top of the statistical uncertainty resulting from the sampling of gauge configurations. However, those results are obtained at an unphysical pion mass of 358 MeV without an attempt at extrapolating to the physical pion mass. Likewise, as we will discuss in detail further in the document, we observe clear signs of systematic effects, either attributable to the lack of continuum limit (discretization effects) or the uncertainty in the light-cone limit of our space-like matrix elements (higher-twist effects). With these limitations in mind, we hope that the results, and in particular the novel characterization of the skewness-dependent moments $A_{3,2},B_{3,2}$ and $A_{4,2},B_{4,2}$ can provide insights for theoretical models of GPDs and extractions from the data.
@@ 2.1 Building light-cone GPDs from space-like matrix elements
where we have used that $\sigma^{\mu\nu}z_{\mu}z_{\nu}=0$. One can immediately identify that $\mathcal{A}_{3}$ disappears in the contraction and $\mathcal{A}_{2,8}$ are $\mathcal{O}(z^{2})$ contaminations. If we introduce Ioffe-time GPDs as:
@@
where scale dependence from regulating the collinear divergences at the $z^{2}\to 0$ limit has been suppressed. We have used the Gordon identity to distribute the contribution of $\mathcal{A}_{5}$ between $H$ and $E$. This relation is equivalent to the one quoted in Bhattacharya et al. (2022) up to choices of conventions. The discussion in Radyushkin (2024) using the DD formalism Radyushkin (1996a, b, 1997) demonstrates that the $\mathcal{A}_{5}$ amplitude is sensitive to the $D$-term in the limit $z^{2}\rightarrow 0$, while the other amplitudes $\mathcal{A}_{1,4,6,7}$ build up the double distributions in a similar fashion as here. The right-hand sides of Eqs. (15) and (16) are very advantageous since they can be evaluated immediately at any value of $z^{2}$ with the assurance that they converge towards the correct light-cone limit for arbitrarily small values of $z^{2}$. However, there is clearly no unique way to construct a Lorentz invariant object at $z^{2}<0$ which converges analytically towards the correct light-cone limit. In the absence of a $D$-term in the non-singlet GPDs that we consider here, we will not use the $\mathcal{A}_{5}$ amplitude at all.
@@ 2.2 Perturbative matching
The matching of unpolarized non-singlet Ioffe-time GPDs in the $\overline{\rm MS}$ scheme to $z^{2}$-dependent pseudo-distributions in coordinate space was first presented at one-loop accuracy in Radyushkin (2019) for the pion’s analog of $\mathcal{A}_{1}$. For clarity, we denote with a bar the $\overline{\rm MS}$ quantity. The matching is given by:
@@ 2.3 Moments of GPDs
In this paper we are concerned with the spin-1/2 unpolarized GPDs, so $K(\alpha,\xi\nu)$ is even in $\xi\nu$, and thus only the $c_{n+1,k}$ where $k$ is even are non-zero. In particular, $c_{n+1,k}$ is always real. The $\xi$ dependence in Eq. (29) is explicitly compatible with the well-known polynomiality property of GPDs.
@@
Then, by isolating the logarithmic part of the kernel, one can introduce an expression resumed at leading-logarithmic accuracy as:
@@
Now we introduce the polynomiality property of GPDs in the absence of a $D$-term because we are dealing with non-singlet distributions:
$\overline{F}_{n}(\xi,t,\mu^{2})=\sum_{\begin{subarray}{c}k=0\\ k\textrm{ even}\end{subarray}}^{n}\xi^{k}A_{n,k}(t,\mu^{2})\,.$ (37)
In matrix form, this reads
$\begin{pmatrix}\overline{F}_{1}(\xi,t,\mu^{2})\\ \overline{F}_{2}(\xi,t,\mu^{2})\\ \overline{F}_{3}(\xi,t,\mu^{2})\\ \overline{F}_{4}(\xi,t,\mu^{2})\\ \cdots\end{pmatrix}=\begin{pmatrix}1&0&0&0&0&0&\\ 0&1&0&0&0&0&\\ 0&0&1&\xi^{2}&0&0&\\ 0&0&0&0&1&\xi^{2}&\\ &&&&&&\ddots\end{pmatrix}\times\begin{pmatrix}A_{1,0}(t,\mu^{2})\\ A_{2,0}(t,\mu^{2})\\ A_{3,0}(t,\mu^{2})\\ A_{3,2}(t,\mu^{2})\\ A_{4,0}(t,\mu^{2})\\ A_{4,2}(t,\mu^{2})\\ \cdots\end{pmatrix}\,,$ (38)
and the final matching is obtained by inserting this expression into Eq. (33), and evaluating the result for a sufficient range of different values of $\xi$ to obtain invertibility (one value of $\xi$ for $F_{1,2}$ and two values for $F_{3,4}$ are enough).
@@ 2.4 Discrete and Lorentz symmetries of Ioffe-time GPDs
Let us first understand the symmetries of the amplitudes $\mathcal{A}_{k}$ entering the Lorentz decomposition of Eq. (10) with respect to $\nu\to-\nu$ and $\xi\to-\xi$. In this section, for brevity we choose a gauge where the link connecting the quark and anti-quark fields is unity. Since it is a gauge invariant quantity the discrete symmetries will not be different in other gauges. The parity (P), charge conjugation (C), and time reversal (T) properties of the Wilson line operator $O^{\mu}(z)=\bar{\psi}(-\frac{z}{2})\gamma^{\mu}\psi(\frac{z}{2})$ in Eq. (3) and its (anti)-hermitian combination $O^{\mu}_{\pm}(z)=\frac{1}{2}\big{(}O^{\mu}(z)\pm O^{\mu}(-z)\big{)}$ are given by:
@@
Finally, let us observe how these discrete symmetries interact with the Lorentz symmetry, which materializes through the polynomiality property, Eq. (37). The feature of polynomiality is an intricate relationship in the $\xi$ dependence of the moments of GPDs. It arises naturally in the double distribution representation, which describes GPDs in $x$-space as (except for the $D$-term which is irrelevant for the non-singlet GPDs we are studying here):
@@ 3 Numerical Implementation
The distillation framework affords two particular advantages over other methods generally employed. The first is that it enables a momentum projection to be performed at each point in our two- and three-point functions. Thus it provides a far better sampling of the gauge configurations. Second, the factorization of correlation functions into the convolution of the perambulators describing the quark propagation and the elementals representing the interpolating operators admits the use of a variational method; we defer the use of this for a later work.
@@ 3.1 Interpolator Construction
Mass eigenstates in the continuum are labelled according to angular momentum and parity quantum numbers $J^{P}$, each of which corresponds to an irreducible representation (irrep) of the continuous rotation group $O\left(3\right)$. Fermionic and bosonic states are subsequently described by double and single-valued representations, respectively. The projection of $J$ along a standard axis, typically taken to be $J_{z}$, provides a further quantum number and is used to label the rows of the representation.
@@ 3.3 From the matrix element to the amplitudes
If $z\neq 0$ and $\vec{q}\cdot\hat{z}=0$, then we observe that, due to our choice of $z^{\mu}=(0,0,0,z_{3})$, only the $M^{3}(p_{f},p_{i},z)$ matrix elements are sensitive to $\mathcal{A}_{2}$ and $\mathcal{A}_{8}$. In practice, this means that the $M^{3}$ matrix elements are entirely and exclusively used to characterize those two amplitudes. Although it has been argued that using $M^{3}$ introduces a finite mixing Constantinou and Panagopoulos (2017), using it does not change the extraction of any other amplitude, and therefore does not impact our light-cone GPDs, which do not involve either $\mathcal{A}_{2}$ or $\mathcal{A}_{8}$.
@@
An alternative possibility to perform the extraction is to remove or average directly the trivial duplicate constraints that the zero singular values represent and form a true matrix inverse. The solution of this linear inverse would exactly reproduce the analog of the analytic formulas presented in Bhattacharya et al. (2022) to extract the amplitudes from the matrix elements. In our calculation, the spinors required to calculate $\langle\langle\Gamma\rangle\rangle_{\mu_{f}\mu_{i}}$ are complicated functions of the momenta, depending upon its magnitude, the specific finite volume irreducible representation, and choices of convention. With many momenta combinations, solving the analytic formulas would require a numerical algebra implementation for practical use. The approach outlined in this section replaces the step of re-deriving these analytic solutions for varying considerations such as off-axis separations or new momentum frames with a generic linear algebra problem.
@@ 4 Analysis of correlation functions
The additional factor of $\sqrt{2}$ is due to the normalization of the isovector vector current.444 The electromagnetic current in the light-quark sector reads $\mathcal{J}^{\mu}=\frac{2}{3}\overline{u}\gamma^{\mu}u-\frac{1}{3}\overline{d}\gamma^{\mu}d$, which can be expressed in terms of isovector $\rho^{\mu}$ and isoscalar $\omega^{\mu}$ components according to $\mathcal{J}^{\mu}=\frac{\rho^{\mu}}{\sqrt{2}}+\frac{\omega^{\mu}}{3\sqrt{2}}$, where $\rho^{\mu}=\left(\overline{u}\gamma^{\mu}u-\overline{d}\gamma^{\mu}d\right)/\sqrt{2}$ and $\omega^{\mu}=\left(\overline{u}\gamma^{\mu}u+\overline{d}\gamma^{\mu}d\right)/\sqrt{2}$. We use the current $\rho^{\mu}$ to probe the isovector flavor structure of the off-forward nucleon matrix elements. We form $R^{\mu}\left(\vec{p}_{f},\vec{p}_{i},z;T,\tau\right)$ by treating both three-point and all two-point correlators as complex-valued functions, as the construction of the subduced helicity operators we use as interpolators, detailed in section 3.1, involves a delicate interplay of complex phases. A depiction of the ratio for a specific choice of $\vec{p}_{f}$ and $\vec{p}_{i}$ is given on the left panel of Fig. 5.
@@
Figure 5: (left) Ratio $R^{0}(\vec{p}_{f},\vec{p}_{i},z=a;T,\tau)$ for $\vec{p}_{f}=(2,1,0)$ and $\vec{p}_{i}=(0,1,2)$ (lattice units) with proton rows $1,1$ for various values of $T$ and $\tau$. We show the result of the matrix element fit using either the summation method or an exponential fit with one excited state, with the exclusion of all $\tau$ and $T-\tau<3$. The data at $T=14$ is dotted for legibility purposes. – (right) Impact of the exclusion of short-time separations below $t_{s}$ on the extraction of the bare matrix element. We show the result of a ground state exponential fit (blue), exponential fit with one excited state (red), and ratio method (green). The standard result in this analysis will be using both the red point at $t_{s}=3$ and at $t_{s}=4$ to evaluate excited state uncertainty.
Figure 6: Dispersion relation $E_{0}\left(|\vec{p}|\right)$ obtained from non-linear fits of the ground state and first excited state on the two-point correlation functions. The dotted line is the continuum dispersion relation.
@@
We first perform a non-linear fit of the ground state and first excited state on the two-point correlation functions, excluding small source-sink separations. For each value of $|\vec{p}|$, we use several rotationally equivalent two-point functions to increase the robustness of our fit. As a result, we obtain for each jack-knife sample the fitted values of $\mathcal{Z}_{n}\left(\vec{p}\right)$ and $E_{n}\left(\vec{p}\right)$ for $n=0$ and $1$. The dispersion relation of the ground state is shown in Fig. 6.

### v2

@@ 1 Introduction
GPDs provide therefore an ideal setup where, at least for the time-being, only lattice QCD can provide systematically controlled information. Besides DVCS and DVMP, an increasing number of exclusive processes with enhanced sensitivity to GPDs are being studied, but their experimental exploration is only at its beginning: double DVCS Guidal and Vanderhaeghen (2003); Belitsky and Mueller (2003), where a recent study of measurability looks promising Deja et al. (2023), di-photon production Pedrak et al. (2020); Grocholski et al. (2022), photon meson pair production Boussarie et al. (2017); Duplančić et al. (2018), or the general category of single diffractive hard exclusive processes Qiu and Yu (2022, 2023a). The maturity of this new phenomenology of GPDs in the perspective of the EIC therefore complements the progress of the extraction of GPDs on the lattice. Preliminary joint lattice-experimental phenomenology of GPDs are already being explored e.g. Guo et al. (2022, 2023); Riberdy et al. (2024).
@@
For the reader interested in the phenomenology of GPDs beyond the framework of lattice calculations, we summarize in Fig. 2 the main physics results of this study, in the form of moments of the isovector GPDs $H^{u-d}$ and $E^{u-d}$:
$\int_{-1}^{1}\mathrm{d}x\,x^{n-1}\begin{pmatrix}H^{u-d}\\ E^{u-d}\end{pmatrix}(x,\xi,t)=\sum_{k=0\textrm{ even}}^{n-1}\begin{pmatrix}A_{n,k}(t)\\ B_{n,k}(t)\end{pmatrix}\xi^{k}\pm\textrm{mod}(n+1,2)\xi^{n}C_{n}(t)\,,$ (1)
where $C_{n}(t)$ are moments of the isovector $D$-term with $+$ for the $H$ GPD and -- for the $E$ GPD. For brevity, the results are quoted at the scale $\mu=2$ GeV in the form of a dipole fit with the value at $t=0$ and a dipole mass $\Lambda_{n,k}$ following the example:
@@
More sophisticated treatment of the data is presented starting from section 5 of this document, in particular using a more flexible parametrization of the $t$-dependence. The quoted uncertainty is constructed by merging the 68% confidence intervals of two different extractions to include an evaluation of excited state contamination. This systematic uncertainty is mostly responsible for the asymmetric errors. A slight asymmetry also arises for statistical reasons for the most poorly determined moments (notably $A_{4,2}$ and $B_{4,2}$) due to the non-linearity of the dipole fit.
These results are obtained at an unphysical pion mass of 358 MeV without an attempt at extrapolating to the physical pion mass. Likewise, as we will discuss in detail further in the document, we observe clear signs of systematic effects, either attributable to the lack of continuum limit (discretization effects) or the uncertainty in the light-cone limit of our space-like matrix elements (higher-twist effects). We use data with separations up to $z=0.6$ fm, which presents some challenges with respect to the procedure of perturbative matching as discussed later. With these limitations in mind, we hope that the results, and in particular the novel characterization of the skewness-dependent moments $A_{3,2},B_{3,2}$ and $A_{4,2},B_{4,2}$ can provide insights for theoretical models of GPDs and extractions from the data.
@@ 2.1 Building light-cone GPDs from space-like matrix elements
where we have used that $\sigma^{\mu\nu}z_{\mu}z_{\nu}=0$. One can immediately identify that $\mathcal{A}_{3}$ disappears in the contraction and $\mathcal{A}_{2,8}$ are $\mathcal{O}(z^{2})$ contaminations since in QCD $\mathcal{A}_{2,8}$ can only have logarithmic colinear divergences in the limit of $z^{2}\rightarrow 0$. If we introduce Ioffe-time GPDs as:
@@
where scale dependence from regulating the collinear divergences at the $z^{2}\to 0$ limit has been suppressed. We have used the Gordon identity to distribute the contribution of $\mathcal{A}_{5}$ between $H$ and $E$. This relation is equivalent to the one quoted in Bhattacharya et al. (2022) up to choices of conventions. The discussion in Radyushkin (2024) using the DD formalism Radyushkin (1996a, b, 1997) demonstrates that the $\mathcal{A}_{5}$ amplitude is sensitive to the $D$-term in the limit $z^{2}\rightarrow 0$, while the other amplitudes $\mathcal{A}_{1,4,6,7}$ build up the double distributions in a similar fashion as here. The right-hand sides of Eqs. (15) and (16) are very advantageous since they can be evaluated immediately at any value of $z^{2}$ with the assurance that they converge towards the correct light-cone limit for arbitrarily small values of $z^{2}$. However, there is clearly no unique way to construct a Lorentz invariant object at $z^{2}<0$ which converges analytically towards the correct light-cone limit.
@@ 2.2 Perturbative matching
Extending the pseudo-PDF’s one-loop matching Radyushkin (2018); Zhang et al. (2018); Izubuchi et al. (2018), the matching of unpolarized non-singlet Ioffe-time GPDs in the $\overline{\rm MS}$ scheme to $z^{2}$-dependent pseudo-distributions in coordinate space was first presented at one-loop accuracy in Radyushkin (2019) for the pion’s analog of $\mathcal{A}_{1}$. For clarity, we denote with a bar the $\overline{\rm MS}$ quantity. The matching is given by:
@@ 2.3 Moments of GPDs
In this paper we are concerned with the spin-1/2 unpolarized GPDs, so $K(\alpha,\xi\nu)$ is even in $\xi\nu$, and thus only the $c_{n+1,k}$ where $k$ is even are non-zero. In particular, $c_{n+1,k}$ is always real.
@@
Then, by isolating the logarithmic part of the kernel, one can introduce an expression resumed at leading-logarithmic accuracy which generalizes the well-known PDF case (for the LL resummation applied to the pseudo-PDF matching, see Gao et al. (2021)) by using the appropriate set of moments diagonalizing the GPD evolution Ohrndorf (1982); Chase (1980):
@@
Now we introduce the polynomiality property of GPDs. It is useful here to separate in the analysis the $D$-term from the rest of the GPD. As we have already mentioned, the $D$-term is contained in a separate Lorentz amplitude $\mathcal{A}_{5}$ Radyushkin (2024), while the rest of the $H$ and $E$ GPDs is constructed out of a combination of the $\mathcal{A}_{1,4,6,7}$ amplitudes. Furthermore, the matching does not mix the $D$-term with the rest of the GPD. To see it, we notice that the polynomiality property writes:
$\overline{F}_{n+1}(\xi,t,\mu^{2})=\sum_{\begin{subarray}{c}k=0\\ k\textrm{ even}\end{subarray}}^{n}\xi^{k}A_{n+1,k}(t,\mu^{2})+\textrm{mod}(n,2)\xi^{n+1}C_{n+1}(t,\mu^{2})\,.$ (37)
where
[…]
are the moments of the $D$-term, while the $A_{n,k}$ do not depend on the $D$-term. Then, the matching formula of Eq. (29) yields:
$\sum_{k=0}^{n}\bigg{[}\sum_{\begin{subarray}{c}i=0\\ i\textrm{ even}\end{subarray}}^{k}\xi^{i}A_{k+1,i}(t,\mu^{2})+\textrm{mod}(k,2)\xi^{k+1}C_{k+1}(t,\mu^{2})\bigg{]}\xi^{n-k}c_{n+1,n-k}$
$\hskip 50.0pt=\sum_{k=0}^{n}\bigg{[}\sum_{\begin{subarray}{c}i=0\\ i\textrm{ even}\end{subarray}}^{k}\xi^{n+i-k}A_{k+1,i}(t,\mu^{2})c_{n+1,n-k}\bigg{]}+\xi^{n+1}\textrm{mod}(n,2)\sum_{k=0}^{n}C_{k+1}(t,\mu^{2})c_{n+1,n-k}\,.$ (39)
The first term only contains powers of $\xi$ up to $\xi^{n}$, while the second term has the same form as a $D$-term. This means that the $z^{2}$ dependent pseudo-GPD obeys a similar polynomiality property as the $\overline{MS}$ GPD, and that the $D$-term of the pseudo-GPD can be matched to the $D$-term of the $\overline{MS}$ GPD independently from the rest of the GPD. It is convenient to write the polynomiality property of the rest of the GPD, noted here as $\overline{F}_{n+1,\textrm{ no D-term}}$ as a simple matrix multiplication:
$\begin{pmatrix}\overline{F}_{1,\textrm{ no D-term}}(\xi,t,\mu^{2})\\ \overline{F}_{2,\textrm{ no D-term}}(\xi,t,\mu^{2})\\ \overline{F}_{3,\textrm{ no D-term}}(\xi,t,\mu^{2})\\ \overline{F}_{4,\textrm{ no D-term}}(\xi,t,\mu^{2})\\ \cdots\end{pmatrix}=\begin{pmatrix}1&0&0&0&0&0&\\ 0&1&0&0&0&0&\\ 0&0&1&\xi^{2}&0&0&\\ 0&0&0&0&1&\xi^{2}&\\ &&&&&&\ddots\end{pmatrix}\times\begin{pmatrix}A_{1,0}(t,\mu^{2})\\ A_{2,0}(t,\mu^{2})\\ A_{3,0}(t,\mu^{2})\\ A_{3,2}(t,\mu^{2})\\ A_{4,0}(t,\mu^{2})\\ A_{4,2}(t,\mu^{2})\\ \cdots\end{pmatrix}\,,$ (40)
and the independent matching of this piece is obtained by inserting this expression into Eq. (33), and evaluating the result for a sufficient range of different values of $\xi$ to obtain invertibility (one value of $\xi$ for $F_{1,2}$ and two values for $F_{3,4}$ are enough).
As for the $D$-term, from:
$\overline{H}(x,\xi,t,\mu^{2})=\overline{H}_{\textrm{ no D-term}}(x,\xi,t,\mu^{2})+\textrm{sgn}(\xi)\overline{D}\left(\frac{x}{\xi},t,\mu^{2}\right)\Theta(|\xi|-|x|)\,,$ (41)
one gets:
[…]
and comparing with Eq. (15) gives that:
[…]
where we have used the oddness in $\alpha$ of the $D$-term which implies that $\mathcal{A}_{5}$ should be purely imaginary in the limit $z^{2}\rightarrow 0$. Deviations from this might arise both because of lattice systematics and higher twist contaminations due to the non-zero value of $z^{2}$. Finally, the introduction of the matching yields:
$\lim_{z^{2}\rightarrow 0}\mathcal{A}_{5}(\nu,\xi,t,z^{2})=i\xi\nu C_{2}(t,\mu^{2})c_{2,0}(\alpha_{s},z^{2}\mu^{2})$
$\hskip 70.0pt-i(\xi\nu)^{3}\bigg{[}C_{4}(t,\mu^{2})c_{4,0}(\alpha_{s},z^{2}\mu^{2})+C_{2}(t,\mu^{2})c_{4,2}(\alpha_{s},z^{2}\mu^{2})\bigg{]}+...$ (45)
@@ 2.4 Discrete and Lorentz symmetries of Ioffe-time GPDs
Let us first understand the symmetries of the amplitudes $\mathcal{A}_{k}$ entering the Lorentz decomposition of Eq. (10) with respect to $\nu\to-\nu$ and $\xi\to-\xi$. In this section, for brevity we choose a gauge where the link connecting the quark and anti-quark fields is unity. Since it is a gauge invariant quantity the discrete symmetries will not be different in other gauges. The parity (P), charge conjugation (C), and time reversal (T) properties of the operator $O^{\mu}(z)=\bar{\psi}(-\frac{z}{2})\gamma^{\mu}\psi(\frac{z}{2})$ in Eq. (3) and its (anti)-hermitian combination $O^{\mu}_{\pm}(z)=\frac{1}{2}\big{(}O^{\mu}(z)\pm O^{\mu}(-z)\big{)}$ are given by:
@@
Finally, let us observe how these discrete symmetries interact with the Lorentz symmetry, which materializes through the polynomiality property, Eq. (37). The feature of polynomiality is an intricate relationship in the $\xi$ dependence of the moments of GPDs. It arises naturally in the double distribution representation, which describes GPDs in $x$-space as (we use only the $D$-term independent part here):
@@ 3 Numerical Implementation
The distillation framework affords two particular advantages over other methods generally employed. The first is that it enables a momentum projection to be performed at each point in our two- and three-point functions. Thus it provides a far better sampling of the gauge configurations. Second, the factorization of correlation functions into the convolution of the perambulators describing the quark propagation and the elementals representing the interpolating operators admits the use of a variational method; we defer the use of this for later work.
@@ 3.1 Interpolator Construction
Mass eigenstates in the continuum are labeled according to angular momentum and parity quantum numbers $J^{P}$, each of which corresponds to an irreducible representation (irrep) of the continuous rotation group $O\left(3\right)$. Fermionic and bosonic states are subsequently described by double and single-valued representations, respectively. The projection of $J$ along a standard axis, typically taken to be $J_{z}$, provides a further quantum number and is used to label the rows of the representation.
@@ 3.3 From the matrix element to the amplitudes
If $z\neq 0$ and $\vec{q}$ is not along the $z$ axis ($q_{1}$ or $q_{2}\neq 0$), then we observe that, due to our choice of $z^{\mu}=(0,0,0,z_{3})$, only the $M^{3}(p_{f},p_{i},z)$ matrix elements are sensitive to $\mathcal{A}_{2}$ and $\mathcal{A}_{8}$. In practice, this means that the $M^{3}$ matrix elements are entirely and exclusively used to characterize those two amplitudes. Although it has been argued that using $M^{3}$ introduces a finite mixing Constantinou and Panagopoulos (2017), using it does not change the extraction of any other amplitude, and therefore does not impact our light-cone GPDs, which do not involve either $\mathcal{A}_{2}$ or $\mathcal{A}_{8}$.
@@
An alternative possibility to perform the extraction is to remove or average directly the trivial duplicate constraints and form a true matrix inverse. The solution of this linear inverse would exactly reproduce the analog of the analytic formulas presented in Bhattacharya et al. (2022) to extract the amplitudes from the matrix elements. In our calculation, the spinors required to calculate $\langle\langle\Gamma\rangle\rangle_{\mu_{f}\mu_{i}}$ are complicated functions of the momenta, depending upon its magnitude, the specific finite volume irreducible representation, and choices of convention. With many momenta combinations, solving the analytic formulas would require a numerical algebra implementation for practical use. The approach outlined in this section replaces the step of re-deriving these analytic solutions for varying considerations such as off-axis separations or new momentum frames with a generic linear algebra problem.
@@ 4 Analysis of correlation functions
The additional factor of $\sqrt{2}$ is due to the normalization of the isovector vector current.444 The electromagnetic current in the light-quark sector reads $\mathcal{J}^{\mu}=\frac{2}{3}\overline{u}\gamma^{\mu}u-\frac{1}{3}\overline{d}\gamma^{\mu}d$, which can be expressed in terms of isovector $\rho^{\mu}$ and isoscalar $\omega^{\mu}$ components according to $\mathcal{J}^{\mu}=\frac{\rho^{\mu}}{\sqrt{2}}+\frac{\omega^{\mu}}{3\sqrt{2}}$, where $\rho^{\mu}=\left(\overline{u}\gamma^{\mu}u-\overline{d}\gamma^{\mu}d\right)/\sqrt{2}$ and $\omega^{\mu}=\left(\overline{u}\gamma^{\mu}u+\overline{d}\gamma^{\mu}d\right)/\sqrt{2}$. We use the current $\rho^{\mu}$ to probe the isovector flavor structure of the off-forward nucleon matrix elements. We form $R^{\mu}\left(\vec{p}_{f},\vec{p}_{i},z;T,\tau\right)$ by treating both three-point and all two-point correlators as complex-valued functions, as the construction of the subduced helicity operators we use as interpolators, detailed in section 3.1, involves a delicate interplay of complex phases. A depiction of the ratio for a specific choice of $\vec{p}_{f}$ and $\vec{p}_{i}$ is given in Fig. 5.
@@
Figure 5: Ratio $R^{0}(\vec{p}_{f},\vec{p}_{i},z=a;T,\tau)$ for $\vec{p}_{f}=(2,1,0)$ and $\vec{p}_{i}=(0,1,2)$ (lattice units) with proton rows $1,1$ for various values of $T$ and $\tau$. We show the result of the matrix element fit using either the summation method or an exponential fit with one excited state, with the exclusion of all $\tau$ and $T-\tau<3$. The data at $T=14$ is dotted for legibility purposes.
Figure 6: (left) Dispersion relation $E_{0}\left(|\vec{p}|\right)$ obtained from non-linear fits of the ground state and first excited state on the two-point correlation functions. The dotted line is the continuum dispersion relation. – (right) Impact of the exclusion of short-time separations below $t_{s}$ on the extraction of the same bare matrix element as in Fig. 5. We show the result of a ground state exponential fit (blue), exponential fit with one excited state (red), and ratio method (green). The standard result in this analysis will be using both the red point at $t_{s}=3$ and at $t_{s}=4$ to evaluate excited state uncertainty.
@@
We first perform a non-linear fit of the ground state and first excited state on the two-point correlation functions, excluding small source-sink separations. For each value of $|\vec{p}|$, we use several rotationally equivalent two-point functions to increase the robustness of our fit. As a result, we obtain for each jack-knife sample the fitted values of $\mathcal{Z}_{n}\left(\vec{p}\right)$ and $E_{n}\left(\vec{p}\right)$ for $n=0$ and $1$. The dispersion relation of the ground state is shown on the left panel of Fig. 6.